# CERN-TH-2025-012, IPPP/25/09, ZU-TH 11/25 NNLOJET: a parton-level event generator for jet cross sections at NNLO QCD accuracy

### v1

@@ 4.5 Photon-plus-jet and di-photon production
The production of isolated photons is a classical collider observable. Its theoretical description must account for the experimental prescription that is used to isolate the photons from hadronic activity in the event. This prescription is typically formulated in terms of an isolation cone around the photon direction, which admits only a limited amount of hadronic energy. All relevant experimental measurements use a fixed-size cone, while theory calculations also consider dynamical isolation cones [110], which gradually reduce the allowed hadronic energy to zero towards the photon direction in the cone centre. Only fixed-size cones admit some amount of collinear photon radiation off QCD partons, thus inducing a dependence on photon fragmentation processes [111]. The NNLOJET implementation of and to NNLO QCD with a dynamical isolation cone is described in [112]. Its extension to fixed-cone isolation is documented in [113], employing an empirical parametrization of the photon fragmentation functions [114]. Di-photon production is currently implemented to NNLO QCD only for dynamical cone isolation [115]. The matrix elements up to NNLO QCD for both processes are taken from [116, 117, 118, 119].
@@ 4.6 Higgs-boson-plus-jet production
The dominant Higgs-boson-production process at the LHC is gluon fusion, which is mediated through a virtual top-quark loop. The NNLOJET implementation of and to NNLO QCD is described in [120]. It is based on an effective field theory (EFT) which couples the Higgs-boson to the gluon field strength, resulting from integrating out the top-quark loop in the infinite-mass limit [121, 122]. The relevant matrix elements up to NNLO are based on [123, 124, 125, 126, 127]. The Higgs boson is considered to be on-shell and a narrow-width approximation is used. Besides the on-shell Higgs production, the Higgs decay modes to [128], and [129] as well as to [130] are implemented, allowing arbitrary fiducial cuts on the Higgs decay products. Heavy-quark mass corrections are implemented at LO QCD for and , allowing for a multiplicative reweighting of the infinite-mass EFT predictions.
@@ 5.3.1 Process block
photon_isolation/lepton_isolation: define the isolation of non-QCD particles, as documented in detail in [113]. It allows the combination of several isolation prescriptions that must be simultaneously fulfilled, using the syntax
@@
hybrid_smooth: hybrid isolation. A variant of the smooth-cone isolation defined in [137]. The following parameters can be specified:
ET_const: ;
@@ 5.3.2 Run block
PDF = <PDF_set>[<member>] [required]: specification of a PDF set according to the naming convention of LHAPDF [67]. For asymmetric colliders, separate PDFs can be specified for each beam by instead providing PDF1 and PDF2 individually. The optional <member> = INT parameter specifies the PDF member. It defaults to 0, i.e. the central member of the PDF set. Note that the PDF is required also for electron-positron collisions, for the running of ;
@@ 5.3.3 Parameters block
The attributes of particle <particle> are set with the syntax <attribute>[<particle>]. <particle> may assume any acronym from Table 1 as its value. The possible attributes are listed in the following, with the default value taken from [139]:
@@
hard_photon_alpha0 = BOOL: switch to use as the electroweak coupling for resolved photons in photon-production processes (default: .false.);
@@
ALPHA0: scheme (override default value with alpha0 = DBLE);
@@ 5.3.6 Scales block
and requires at least one scale specification to be present. The first entry will be considered the central scale that is also used for the Vegas grid adaption. Up to six additional scale specifications can be supplied. A scale specification constitutes a line in the block that specifies the factorization scale (muF) and the renormalization scale (muR)
In case a process that contains a fragmentation function is run, the associated factorization scale () can be set separately using a pair muF = [<sclF>,<sclA>], which otherwise will default to in case only a single scale is specified. The scale <scl> can be set in two ways:
@@ 5.3.8 Single-line Optional Settings
The PDF interpolation option is a single-line setting to switch on a one-dimensional cubic spline interpolation (fixed , interpolation in ), which aims to reduce the number of calls to LHAPDF. By default it is not activated but it can be switched on with the following syntax:
@@ 6.1 Job submission workflow
order = lo|nlo|nlo_only|nnlo(%/key)|(key)nnlo_only: specifies the perturbative order of the calculation to be performed. It is also possible to restrict the calculation to only a specific coefficient, nnlo=lo+nlo_only+nnlo_only. (default: nnlo)
@@
The finalize sub-command facilitates a manual trigger to perform a re-combination of results with changed internal parameters of the merging algorithm. The main purpose of the merging algorithm is to combine the partial results of different parts of the calculation into a final prediction and to mitigate the impact from outliers. Potential issues pertaining to the treatment of outliers, as well as the approach employed in our implementation are presented in Ref. [141]. The underlying algorithm is controlled by a set of parameters that have sensible defaults but can also be overridden by the user. This is useful to check if the combine step is potentially introducing a systematic bias as well as to optimize the combination to obtain the best possible result from the raw data. The merging proceeds in steps with a dynamical termination condition, which are applied for each histogram separately on a bin-by-bin basis:
@@ Acknowledgements
The NNLOJET project has been enabled by substantial support throughout the past decade from the European Research Council (ERC) through grant agreements 340983 (ERC Advanced Grant MC@NNLO) and 101019620 (ERC Advanced Grant TOPUP), from the Swiss National Science Foundation (SNF) under contracts 200020-162487, 200021-172478, 200020-175595, 200021-197130, 200020-204200 and 200021-231259, from the Swiss National Supercomputing Centre (CSCS) under project IDs ETH5f and UZH10, from the Royal Society, from the UK Science and Technology Facilities Council (STFC) under contracts ST/G000905/1, ST/P001246/1, ST/T001011/1 and ST/X000745/1, and from the National Science Foundation of China (NSFC) under contract No.12475085, No.12321005 and No.1223500. MM is supported by a Royal Society Newton International Fellowship (NIF/R1/232539). JP acknowledges Fundação para a Ciência e a Tecnologia (FCT) for its financial support through the programatic funding of R&D units (contract UIDP/50007/2020) and under project CERN/FIS-PAR/0032/2021.

### v2

@@ 4.5 Photon-plus-jet and di-photon production
The production of isolated photons is a classical collider observable. Its theoretical description must account for the experimental prescription that is used to isolate the photons from hadronic activity in the event. This prescription is typically formulated in terms of an isolation cone around the photon direction, which admits only a limited amount of hadronic energy. All relevant experimental measurements use a fixed-size cone, while theory calculations also consider dynamical isolation cones [Frixione:1998jh], which gradually reduce the allowed hadronic energy to zero towards the photon direction in the cone centre. Only fixed-size cones admit some amount of collinear photon radiation off QCD partons, thus inducing a dependence on photon fragmentation processes [Koller:1978kq]. The NNLOJET implementation of and to NNLO QCD with a dynamical isolation cone is described in [Chen:2019zmr]. Its extension to fixed-cone isolation is documented in [Chen:2022gpk], employing an empirical parametrization of the photon fragmentation functions [Bourhis:1997yu]. Di-photon production is currently implemented to NNLO QCD only for dynamical cone isolation [Gehrmann:2020oec]. The matrix elements up to NNLO QCD for both processes are taken from [DelDuca:1999pa, Signer:1995np, Anastasiou:2002zn, Campbell:2014yka].
@@ 4.6 Higgs-boson-plus-jet production
The dominant Higgs-boson-production process at the LHC is gluon fusion, which is mediated through a virtual top-quark loop. The NNLOJET implementation of and to NNLO QCD is described in [Chen:2014gva]. It is based on an effective field theory (EFT) which couples the Higgs-boson to the gluon field strength, resulting from integrating out the top-quark loop in the infinite-mass limit [Shifman:1978zn, Wilczek:1977zn]. The relevant matrix elements up to NNLO are based on [DelDuca:2004wt, Dixon:2009uk, Badger:2009hw, Badger:2009vh, Gehrmann:2011aa]. The Higgs boson is considered to be on-shell and a narrow-width approximation is used. Besides the on-shell Higgs production, the Higgs decay modes to [Chen:2016zka], and [Chen:2019wxf] as well as to [Chen:2021ibm] are implemented, allowing arbitrary fiducial cuts on the Higgs decay products. Heavy-quark mass corrections are implemented at LO QCD for and , allowing for a multiplicative reweighting of the infinite-mass EFT predictions.
@@ 5.3.1 Process block
anti- algorithm [Cacciari:2008gp]: antikt[DBLE];
Cambridge–Aachen algorithm [Wobisch:1998wt] (longitudinally invariant): ca[DBLE];
algorithm [Ellis:1993tq, Catani:1993hr] (longitudinally invariant): kt[DBLE];
Durham algorithm [Catani:1991hj] (for collisions): durham[DBLE];
Jade algorithm [JADE:1986kta] (for collisions): jade[DBLE];
No algorithm required: none;
@@
photon_isolation/lepton_isolation: define the isolation of non-QCD particles, as documented in detail in [Chen:2022gpk]. It allows the combination of several isolation prescriptions that must be simultaneously fulfilled, using the syntax
@@
smooth: smooth-cone isolation [Frixione:1998jh]. The non-QCD particle with transverse momentum is isolated according to a profile function that permits a decreasing amount of hadronic energy within sub-cones of radius and as defined in (9):
hybrid_smooth: hybrid isolation. A variant of the smooth-cone isolation defined in [Siegert:2016bre]. The following parameters can be specified:
ETmax_const: ;
@@
ALEPH: a simple parametrization obtained from a one-parameter fit by ALEPH [ALEPH:1995zdi];
BFG0, BFGPERT: BFG0 [Bourhis:1997yu];
BFG1, BFGI: BFGI [Bourhis:1997yu];
BFG2, BFGII: BFGII [Bourhis:1997yu].
@@ 5.3.2 Run block
PDF = <PDF_set>[<member>] [required]: specification of a PDF set according to the naming convention of LHAPDF [Buckley:2014ana]. For asymmetric colliders, separate PDFs can be specified for each beam by instead providing PDF1 and PDF2 individually. The optional <member> = INT parameter specifies the PDF member. It defaults to 0, i.e. the central member of the PDF set. Note that the PDF is required also for electron-positron collisions, for the running of s;
@@ 5.3.3 Parameters block
The attributes of particle <particle> are set with the syntax <attribute>[<particle>]. <particle> may assume any acronym from Table 1 as its value. The possible attributes are listed in the following, with the default value taken from [ParticleDataGroup:2018ovx]:
@@
hard_photon_alpha0 = BOOL: switch to use 0 as the electroweak coupling for resolved photons in photon-production processes (default: .false.);
@@
ALPHA0: 0 scheme (override default value with alpha0 = DBLE);
@@
IDENTITY (default): diagonal CKM matrix;
FULL: Wolfenstein parameters [ParticleDataGroup:2020ssz];
CABIBBO: Cabibbo mixing angle only [ParticleDataGroup:2020ssz];
MCFM: the default setup of MCFM.
@@ 5.3.6 Scales block
and requires at least one scale specification to be present. The first entry will be considered the central scale that is also used for the Vegas grid adaption. Up to six additional scale specifications can be supplied. A scale specification constitutes a line in the block that specifies the factorization scale F (muF) and the renormalization scale R (muR)
In case a process that contains a fragmentation function is run, the associated factorization scale (A) can be set separately using a pair muF = [<sclF>,<sclA>], which otherwise will default to in case only a single scale is specified. The scale <scl> can be set in two ways:
@@ 5.3.8 Single-line Optional Settings
The PDF interpolation option is a single-line setting to switch on a one-dimensional cubic spline interpolation (fixed , interpolation in F), which aims to reduce the number of calls to LHAPDF. By default it is not activated but it can be switched on with the following syntax:
@@ 6.1 Job submission workflow
order = lo|nlo|nlo_only|nnlo|nnlo_only: specifies the perturbative order of the calculation to be performed. It is also possible to restrict the calculation to only a specific coefficient, nnlo=lo+nlo_only+nnlo_only. (default: nnlo)
@@
The finalize sub-command facilitates a manual trigger to perform a re-combination of results with changed internal parameters of the merging algorithm. The main purpose of the merging algorithm is to combine the partial results of different parts of the calculation into a final prediction and to mitigate the impact from outliers. Potential issues pertaining to the treatment of outliers, as well as the approach employed in our implementation are presented in Ref. [Gehrmann-DeRidder:2016ycd]. The underlying algorithm is controlled by a set of parameters that have sensible defaults but can also be overridden by the user. This is useful to check if the combine step is potentially introducing a systematic bias as well as to optimize the combination to obtain the best possible result from the raw data. The merging proceeds in steps with a dynamical termination condition, which are applied for each histogram separately on a bin-by-bin basis:
@@ Acknowledgements
The NNLOJET project has been enabled by substantial support throughout the past decade from the European Research Council (ERC) through grant agreements 340983 (ERC Advanced Grant MC@NNLO) and 101019620 (ERC Advanced Grant TOPUP), from the Swiss National Science Foundation (SNF) under contracts 200020-162487, 200021-172478, 200020-175595, 200021-197130, 200020-204200 and 200021-231259, from the Swiss National Supercomputing Centre (CSCS) under project IDs ETH5f and UZH10, from the Royal Society, from the UK Science and Technology Facilities Council (STFC) under contracts ST/G000905/1, ST/P001246/1, ST/T001011/1 and ST/X000745/1, and from the National Science Foundation of China (NSFC) under contract No.12475085, No.12321005 and No.1223500. MM is supported by a Royal Society Newton International Fellowship (NIF/R1/232539). JP acknowledges Fundação para a Ciência e a Tecnologia (FCT) for its financial support through the programatic funding of R&D units (contract UIDP/50007/2020) and under project CERN/FIS-PAR/0032/2021. The work of LB, XC, GF, MJ, and KS was supported by the UZH Forschungskredit Candoc and Postdoc schemes.
Appendix A Example: at NNLO
In this Appendix, we highlight an example calculation for the process at NNLO. The full runcard is provided below, implementing the following items:
In the PROCESS block, we specify the process (-boson plus one jet) and the collider (proton–proton collider at . We further specify the jet algorithm used to reconstruct jets (anti- with ). See Sect. 5.3.1.
In the RUN block, we specify the PDF set, here NNPDF3.1. Additionally, we indicate a value for the technical cutoff (equal to the default ) and we turn on the multi-channel adapatation over individual channels (in this case, channels will be grouped such that every phase-space point will evaluate a subset of 5 channels). See Sect. 5.3.2.
In the PARAMETERS block, we select the scheme and set the values for the masses and widths of the and bosons and for the Fermi constant . See Sect. 5.3.3.
In the SELECTORS block, we define the fiducial phase space for the selection of events. In particular, we require leptons to have and . Moreover, their angular distance in the rapidity–azimuth plane should be and they should feature an invariant mass . Jets are required to have and , with . See Sect. 5.3.4.
In the HISTOGRAMS block, we declare some histograms, such as the transverse momentum or the absolute rapidity of the leading- jet. Note in particular some features like: the possibility of specifying uniformly-spaced bins or declaring them with a custom list; the file-name override (ptj1 > ptj1_exc) to avoid name clashes when saving histograms; the histogram selectors to generate multi-differential distributions (e.g. we fill the ptz_y1 histogram with the transverse momentum of the -boson only when ). See Sect. 5.3.5.
In the SCALES block, we specify the set of renormalization and factorization scales to be registered for the calculation: we adopt (with the scalar- sum over leptons and jets) as central scale and we further compute six additional combinations of R and F according to the usual 7-point scale variation prescription to estimate theory uncertanties. See Sect. 5.3.6.
The above runcard can be used as input to the job-submission workflow described in Sect. 6.1. After a run on the CERN lxplus submission system with a total runtime of order 50k CPU hours, one obtains a value of pb for the total NNLO fiducial cross section. Figure 2 shows an example of a differential distribution, namely the transverse momentum of the leading jet. The runcard with all generated data files and plots are included as ancillary files to this publication.
We remind the reader that more example runcards, including reference results, are available on the website https://nnlojet.hepforge.org/examples.html.